1. Measuring employment, unemployment, and labor force participation

Consider a small economy composed of six people: Van, Amy, Carlos, Deborah, Janet, and Felix. Each person's employment status is described in the following table.

Based on the criteria used by the Bureau of Labor Statistics (BLS), identify each person’s status as employed, unemployed, “not in the labor force” (if not in the civilian labor force but still part of the adult population), or “not in the adult population” if not in the civilian adult population.

Person

Status

Van is a 75-year-old professor. He teaches only one or two courses a year, but he's still pursuing an active research agenda.   
Amy is a 27-year-old recent college graduate. She did not work for pay last week, but she had two job interviews.   
Carlos is a 78-year-old retired professor. He enjoys volunteering at the local public library.   
Janet is a 29-year-old who lost her job as an associate producer for a radio station. After spending a few weeks out of work and interviewing for several other positions, she gave up on her job search and decided to go back to grad school. She made that decision a few months ago.   
Felix is a famous novelist. He is spending the summer at his lake house in upstate New York, doing a little writing each day but mostly spending his time gardening and reading.   
Deborah is a 12-year-old student at East Valley Middle School. She babysits her younger brother and does other chores, so her parents give her an allowance of $30 per week.   

Complete the right half of the following equation to reflect the labor force participation rate reported by the BLS.

Labor Force Participation RateLabor Force Participation Rate =  =   

According to this formula, what is the labor force participation rate of this economy of six people?

33.33%

40%

50%

60%

Complete the right half of the following equation to reflect the unemployment rate reported by the BLS.

Unemployment RateUnemployment Rate =  =   

According to this formula, what is the unemployment rate of this economy of six people?

33.33%

40%

50%

60%
Answer #1

Van: Employed

Amy: Unemployed

Carlos: Not in labor force

Janet: Not in labor force

Felix: Employed

Deborah: Not in adult population

Labor force participation rate = Labor Force / Adult population * 100

Labor force = Employed + unemployed = 2+1 = 3

Adult popolation = Employed + unemployed + not in labor force = 2+1+2 = 5

Labor force participation rate = 3/5 * 100 = 60%

Unemployment rate = Unemployed / labor force * 100

Unemployment rate = 1/3 * 100 = 33.33%